# Similarity of Neural Architectures using Adversarial Attack Transferability

### v3

@@ G.2 Teacher accuracy and distillation performance
Figure G.1: Teacher network accuracy and distillation performance. The trend line and its 90% confidence interval are shown in the blue line and gray band. There is no significant correlation between teacher accuracy and distillation performance.
@@
The similarity between teacher and student networks may not be the only factor contributing to distillation. For example, a stronger teacher can lead to better distillation performance [115]. In Fig. G.1, we observe that if the range of the teacher accuracy is not significantly large enough (e.g., between 79.5 and 82.5), then the correlation between teacher network accuracy and distillation performance is not significant; 0.049 Pearson correlation coefficient with 0.82 p-value. In this case, we can conclude that the teacher and student networks similarity contributes more to the distillation performance than the teacher performance.
Appendix H Limtations and discussions
Figure H.1: Similarity scores using two-side attack transferability vs. one-side attack transferability. The trend line and its 90% confidence interval are shown in the blue line and gray band. We show the relationship between similarity scores using two-side and one-side attack transferability.
Similarity scores for adversarially trained networks.
Because our method is based on an adversarial attack, it could be affected by an adversarial attack-aware learning method. We first note that our analysis is invariant to the adversarial training; as shown by Tsipras et al. [100], adversarial robustness and accuracy are in a trade-off, and there is no ImageNet model with accuracy aligned with our target models yet. Also, our model similarity is based on the difference in the change of the model predictions. As shown by Tsipras et al. [100] and Ilyas et al. [49], adversarial training will lead to a different decision boundary. Therefore, we think that although our score can be affected by adversarial training, our score is still valid for adversarially trained models because an adversarially trained model is actually dissimilar to other models in terms of the decision boundary and representation learning.
An efficient approximation for a new model.
@@ Appendix H Limtations and discussions
We can use our toolbox for designing a new model; we can measure the similarity between a novel network and existing $N$ architectures; a novel network can be assigned to clusters (Tab. 2) to understand how it works. However, it requires generating adversarial samples for all $N$+1 models (e.g., 70 in our case), which is computationally inefficient. Instead, we propose the approximation of Eq. (1) of the main paper by omitting to compute the accuracy of the novel network on the adversarial samples of the existing networks. It will break the symmetricity of our score, but we found that the approximated score and the original score have high similarity (0.82 Pearson correlation coefficient with almost 0 p-value) as shown in Fig. H.1).
@@
As an example, we tested Gluon-ResNeXt-50 [37], which has the same architecture with ResNeXt-50 [110] in 69 models, and the distilled version of DeiT-S model [96]. Because the difference is not significant as we observed in Tab. 2 and Fig. 4 (b) of the main paper, we expect that Gluon-ResNeXt-50 is assigned to the same cluster with ResNeXt-50, and distilled DeiT-S is assigned to the same cluster with DeiT-S. We observe that, as we expected, each network is assigned to the desired cluster. Therefore, we suggest using our efficient approximation for analyzing a novel network with our analysis toolbox.
Other possible applications.
We can use our similarity to develop a novel neural architecture or training technique by measuring how the proposed method can learn different representations compared to existing methods. Our similarity also can be used for analyzing the hyper-parameter sensitivity. Instead of reporting the performance changes by different hyper-parameters, one can report the average similarity of the trained models with different hyper-parameters. If the average similarity is significantly low, then we can argue that the method is sensitive to hyper-parameter selection. We expect this scenario is especially useful for the task, which is sensitive to the hyper-parameter selection [16, 36]. Finally, our similarity can be used for the model selection criteria for selecting diverse models, e.g., constructing a new dataset with diverse pre-trained models for annotation candidates [59, 20].

### v4

@@ 0.F.1 Teacher Accuracy and Distillation Performance
Figure 0.F.1: Teacher network accuracy and distillation performance. There is no significant correlation between teacher accuracy and distillation performance.
@@
The similarity between teacher and student networks may not be the only factor contributing to distillation. For example, a stronger teacher can lead to better distillation performance [123]. In Fig. 0.F.1, we observe that if the range of the teacher accuracy is not significantly large enough (e.g., between 79.5 and 82.5), then the correlation between teacher network accuracy and distillation performance is not significant; 0.049 Pearson correlation coefficient with 0.82 p-value. In this case, we can conclude that the teacher and student networks similarity contributes more to the distillation performance than the teacher performance.
Appendix 0.G Limtations and Discussions
0.G.1 Efficient Approximation of SAT for a Novel Model
@@
We can use our toolbox for designing a new model; we can measure SAT between a novel network and existing $N$ architectures; a novel network can be assigned to clusters (Tab. 1) to understand how it works. However, it requires generating adversarial samples for all $N+1$ models (e.g., 70 in our case), which is computationally inefficient. Instead, we propose the approximation of Eq. 1 by omitting to compute the accuracy of the novel network on the adversarial samples of the existing networks. It will break the symmetricity of SAT, but we found that the approximated score and the original score have high similarity – 0.82 Pearson coefficient with almost 0 p-value – as shown in Fig. 1(a).
(a) Approximated (one-side) SAT vs. original (two-side).
(b) SAT vs. The number of same misclassification.
Figure 0.G.1: Additional Analysis for SAT.
@@
As an example, we tested Gluon-ResNeXt-50 [39] and the distilled version of DeiT-S [100]. As observed in Tab. 2 and Fig. 6, models with the same architecture have high similarity compared to models with different architectures; hence, we expect that Gluon-ResNeXt-50 is assigned to the same cluster with ResNeXt-50, and distilled DeiT-S is assigned to the same cluster with DeiT-S. As we expected, each network is assigned to the desired cluster. Therefore, we suggest using our efficient approximation for analyzing a novel network with our analysis toolbox.
0.G.2 Adversarial Attack Transferability and Direction of Missclassification
Waseda et al. [113] showed that adversarial attack transferability is highly related to the direction of the misclassification. We examine if SAT is related to the misclassification. Fig. 1(b) shows the relationship between SAT and the number of the same misclassification by the attack. We observe that they are highly correlated, i.e., we confirmed that SAT is also related to the misclassification.
0.G.3 Change of SAT During Training
We check the adversarial attack transferability between the fully trained model and less trained models on CIFAR-10 with 180 training epochs. A model trained with only 20 epochs shows high similarity over different initializations (4.23 in Tab. 2 of the main paper). Note that SAT considers models having similar clean accuracy; namely, there is room to explore this further in future work.
Table 0.G.1: Change of SAT between fully-trained model (epoch 180) and models on various epochs.
Epoch 0 20 40 60 80 100 120 140 160 180
SAT 2.84 4.46 4.56 4.58 4.59 4.59 4.60 4.60 4.60 4.60
0.G.4 More Possible Applications Requiring Diverse Models
[…]